Purchase-Related Transactions Warwick’s Co., a women's clothing store, purchased $21,000 of merchandise from a supplier on account, terms FOB destination, 2/10, n/30. Warwick’s returned $3,400 of the merchandise, receiving a credit memo, and then paid the amount due within the discount period. a. Journalize Warwick’s Co.'s entry to record the purchase. Merchandise Inventory 73,500 Accounts Payable 73,500 Feedback a. A purchase increases Merchandise Inventory and Accounts Payable. b. Journalize Warwick’s Co.'s entry to record the merchandise return. Accounts Payable 8,820 Merchandise Inventory 8,820 Feedback b. A return reduces Merchandise Inventory and Accounts Payable in the purchaser's records. c. Journalize Warwick’s Co.'s entry to record the payment. Accounts Payable 64,680 Cash 64,680
Answer #1
a Merchandise Inventory 20580 =21000*(1-2%)
      Accounts Payable 20580
b Accounts Payable 3332 =3400*(1-2%)
      Merchandise Inventory 3332
c Accounts Payable 17248 =20580-3332
      Cash 17248
Note: The solution is based on Net method